OpenStudy (anonymous):

Given that the line Ax+By+10=0 is parallel to the line 3x+y=7 and meets the line x+4=7 on the x-axis, find the value of A and B..??

OpenStudy (anonymous):

How do i do this??

OpenStudy (anonymous):

@mathmate @jim_thompson5910

OpenStudy (anonymous):

do u know??

OpenStudy (jackthegreatest):

maybe...

OpenStudy (jackthegreatest):

since the lines r parallel, the slope is -3

OpenStudy (jackthegreatest):

so A is 3

OpenStudy (anonymous):

wait how A=3

OpenStudy (jackthegreatest):

because the line is parallel to 3x + y= 7 parallel lines have the same slope

OpenStudy (jackthegreatest):

B is 1

OpenStudy (jackthegreatest):

so a=3 and b = 1

OpenStudy (jackthegreatest):

there r ur answers : )

OpenStudy (anonymous):

r u sure??

OpenStudy (jackthegreatest):

yes i am sure, of course i am

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

jackthegreatest is correct

OpenStudy (anonymous):

so we take the slope and use it for A?

jimthompson5910 (jim_thompson5910):

Anything parallel to Ax+By+10=0 will be of the form Ax+By = C where C is some fixed number Using numbers anything parallel to 3x+y=7 will be in the form 3x+y = C

OpenStudy (anonymous):

ok and since y=1 we use that for B?

jimthompson5910 (jim_thompson5910):

the coefficient for y is 1, yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so if we had 4x+5y=7 then A=4 and B=5?

jimthompson5910 (jim_thompson5910):

yeah because the left side of a standard form equation stays the same when you have parallel lines

OpenStudy (anonymous):

ok thank u both.. @jackthegreatest and @jim_thompson5910

jimthompson5910 (jim_thompson5910):

np

OpenStudy (jackthegreatest):

np